OpenStudy (anonymous):

express in a single logarithm 4 log6 4+ 1/4 log6 (x-3)-1/2 log6 x

OpenStudy (anonymous):

hm does it look like this...

OpenStudy (anonymous):

\[(log\ a)+(log\ b) = log(a*b) \] \[3(log\ a) = (log\ a) + (log\ a) + (log\ a) = log(a*a*a) = log(a^3)\] See if that doesn't help with the simplification.